Use the given information to find the

​p-value.

​Also, use a 0.05 significance level and state the conclusion about the null hypothesis​ (reject the null hypothesis or fail to reject the null​ hypothesis).With

Upper H 1H1​:

pgreater than>​0.554,

the test statistic is

zequals=1.34.

A.

​0.0901; reject the null hypothesis

B.

​0.0901; fail to reject the null hypothesis

C.

​0.9099; fail to reject the null hypothesis

D.

​0.1802; reject the null hypothesis
Answer #1

Solution:

This is a right tailed test,

P-value = P( Z > z )

= 1 - P(Z < 1.34 )

= 1 - 0.9099

= 0.0901

Therefore, p-value > 0.05 , Fail to reject the null hypothesis.

C)

​0.9099; fail to reject the null hypothesis.